### v3

# One-Versus-Others Attention: Scalable Multimodal Integration for Clinical Data

Abstract
Multimodal models have become increasingly important as they surpass single-modality approaches on diverse tasks ranging from question-answering to autonomous driving. Despite the importance of multimodal learning, existing efforts focus on vision-language applications, where the number of modalities rarely exceeds four (images, text, audio, video). However, data in other domains, such as healthcare, may include many more modalities like X-rays, PET scans, MRIs, genetic screening, genomic data, and clinical notes, creating a need for both efficient and accurate data integration. Many state-of-the-art multimodal models rely on cross-attention or self-attention for effective data integration, which do not scale well for applications with more than two modalities. The complexity per layer of computing attention in either paradigm is, at best, quadratic with respect to the number of modalities, posing a computational bottleneck that impedes broad adoption. To address this, we propose a new attention mechanism, One-Versus-Others (OvO) attention, that scales linearly with the number of modalities, thus offering a significant reduction in computational complexity compared to existing multimodal attention methods. Using four clinical datasets with diverse modalities, we show that our method decreases computation costs while maintaining or increasing performance compared to popular integration techniques. Across all clinical datasets, OvO reduced the number of required floating point operations (FLOPs) by at least 91.98%, demonstrating its significant impact on efficiency and enabling wider adaptation. 111Code is available at https://anonymous.4open.science/r/OvO-13BD/.
1 Introduction
Multimodal learning has emerged as a promising approach, which enables joint learning from multiple modalities of data (e.g., text and images). Combining different modalities allows for a more comprehensive and accurate understanding of tasks such as clinical decision support (Ming et al., 2022; Golovanevsky et al., 2022; Hayat et al., 2022), image and video captioning (Yu et al., 2019; Seo et al., 2022), audio-visual speech recognition (Sterpu et al., 2018), and sentiment analysis (Poria et al., 2018). Multimodal learning has been explored through various methods in machine learning and deep learning. While feature-level integration was mostly used in more traditional machine learning algorithms, Neural Networks have allowed for the intermediate fusion of modalities at any layer and late fusion at the decision-making stage. However, these fusion paradigms lack a key component - capturing explicit interaction between modalities. For example, in detecting Alzheimer’s Disease, genetic features help reinforce and ground the clinical information and thus lead to more robust decision-making (Golovanevsky et al., 2022). Such relevant interactions can be captured through the attention mechanism. Popular multimodal models, such as LXMERT (Tan and Bansal, 2019) and ViLBERT (Lu et al., 2019), use a fusion method that captures interactions between modalities using cross-attention. On the other hand, models such as VisualBERT (Li et al., 2019) and VL-BERT (Su et al., 2019) use early fusion, where vision and language inputs are concatenated early to learn multimodal through self-attention.
However, both self-attention and cross-attention grow quadratically in computational burden with the number of modalities, posing a scalability challenge. While in popular vision-language integration tasks, the number of modalities rarely exceeds four (images, text, audio, video), a significant bottleneck can arise in other domains. The healthcare domain exemplifies this issue, as a single task may involve integrating data from complex and rich sources spanning multiple modalities from radiology, pathology, genomics, genetics, and clinical data. Therefore, with the influx of many modalities, the use of cross-attention or self-attention will remain limited in the clinical domain as their computational demands escalate even further. To address this gap, we propose a new attention mechanism, One-Versus-Others (OvO) attention. OvO attention is calculated by comparing one modality against a combined representation of all other modalities (hence the name, One-Versus-Others). Our approach significantly reduces computational complexity as it grows linearly with the number of modalities (see Section 3.3). Figure 1 sketches a four-modality example to demonstrate the difference between our approach (scales linearly) and self-attention/cross-attention (scales quadratically). OvO is a general attention scheme that can be integrated into existing multimodal architectures instead of cross-attention or self-attention.
We first, present a complexity analysis and demonstrate its validity through a simulated dataset. Our simulation results show scalability gains in an extreme multimodal setting (20 modalities). Next, we use four diverse clinical datasets that vary in modalities, encoder types (pre-trained and not), number of samples, and tasks to show our model’s improved scalability in different clinical multimodal settings. Our results demonstrate that our method dramatically decreases computation costs (offering at minimum a 91.98% reduction in computations), compared to self-attention and cross-attention while maintaining or even exceeding performance. Finally, we demonstrate the generalizability of our method on an additional non-clinical dataset, further underscoring OvO’s broad applicability.
Overall, OvO is a novel attention scheme for multimodal integration that is domain-agnostic and scales linearly to the number of modalities. Importantly, it enables the practical application of deep learning models in healthcare, where computational efficiency and accuracy are vital for deployment.
2 Related work
Multimodal attention-based models are increasingly pivotal in clinical decision support systems, paralleling their widespread use in vision-language applications. In the medical domain, these models have shown remarkable utility in diverse scenarios, such as cancer classification (LI et al., 2021), biomarker discovery (Braman et al., 2021; Ilyin et al., 2004), prognosis prediction (Schulz et al., 2021; Silva and Rohr, 2020), and more. These applications highlight the versatility and potential of multimodal learning in handling complex and rich medical data. The attention mechanism serves as a core component in these models. Attention measures the similarity among individual representations, like word embedding vectors or, in the multimodal scenario, modality-specific embeddings. Each input embedding can assume one of three roles: (1) Query (), representing the current focus of attention when compared against other input embeddings; (2) Key (), signifying an input embedding being compared to the Query; and (3) Value (), which contributes to computing the output for the Query.
Commonly, the representations from each modality in the multimodal models are passed through one of two paradigms - early fusion followed by self-attention or fusion through cross-attention. The early fusion group (e.g., Transformer-based models like Uniter (Chen et al., 2020), VisualBERT (Li et al., 2019), Vl-BERT (Su et al., 2019)) concatenates the visual embeddings and the textual embeddings as a single input, before passing the inputs through attention (see Figure 1 (a)). Given modalities and , queries (), keys (), and values () are computed from their concatenated sequence (e.g., ). The final output from a standard Transformer block is denoted by , Equation 1 shows the early fusion paradigm.
The cross-attention scheme (used in Transformer-based models like ViLBERT (Lu et al., 2019), LXMERT (Tan and Bansal, 2019), ActBERT (Zhu and Yang, 2020), MulT (Tsai et al., 2019), etc.) inputs each modality into its own Transformer, the outputs of which are fed to a cross-modal Transformer (see Figure 1 (b)). For such models, the cross-modal interactions are captured in a pairwise manner through cross-attention, where queries (), keys (), and values () are computed from the modality inputs ( and ), and then the keys and values from each modality are fed to the multi-headed attention block of the other modality. The output, , is shown in Equation 2.
While the early fusion and cross-attention paradigms could be extended to three modalities, seen in TriBERT (Rahman et al., 2021) and VATT (Akbari et al., 2021), these models face scalability challenges for more than three modalities. Cross-attention methods can leverage joint representations formed from cross-attention but do not scale well to larger numbers of modalities as they are computed in a pairwise fashion. Thus, if there are modalities, computing pairwise fusion between each pair will result in matrix computations. Moreover, attention is not a symmetric calculation, which means that most commonly, it is computed bi-directionally (e.g., image to text and text to image), leading to an even greater computational burden. Early fusion involves the concatenation of modalities before the Transformer layer, which similarly does not scale well with the number of modalities. Self-attention is quadratic with respect to sequence length (Vaswani et al., 2017), and since early fusion methods concatenate inputs before attention, the computational complexity will increase quadratically as the number of modalities increases (see Section 3.3). Furthermore, concatenation is not order invariant, making the ordering of modalities an important consideration, potentially requiring similar bi-directional computations as cross-attention. Our integration method, OvO, addresses the limitations mentioned above in a scalable and domain-agnostic manner.
3 Methods
3.1 One-Versus-Others (OvO) attention
We propose a new attention mechanism, One-Versus-Others (OvO) Attention, which grows linearly with number of modalities rather than quadratically, as would be the case for cross-attention or self-attention (see Section A). OvO computes attention between one modality at a time with respect to all other modalities. Given modality , which is an embedding obtained from a dedicated encoder, where is the number of modalities and , OvO takes in one modality and computes the dot product against all the other modalities with a weight matrix . is a learnable parameter that is shared across all modalities (see Figure 1 (c)) and can learn interactions throughout training. The similarity score function, representing the degree of alignment between the chosen modality and others, calculated for modality with respect to a set of other modalities () is shown in Equation 3. The context vector in OvO for modality , which is a combined representation of information from the other modalities, is shown in Equation 4:
We chose to sum over the “other” modalities instead of concatenation for two reasons: (1) the concatenation vector will continue to increase in length with the number of modalities, which will result in a less scalable framework; (2) concatenation is not invariant to the order of modalities, which could affect the model prediction, whereas a sum provides position invariance.
Note that for modalities (, ), the similarity score function simplifies to that of general attention (Luong et al., 2015):
3.2 Multi-headed OvO Attention
We extend OvO attention to the multi-headed attention framework to directly compare with early fusion through self-attention and pairwise cross-attention. Multi-headed attention allows the model to attend to the input embeddings in different ways simultaneously. This is achieved by splitting the input embeddings into multiple linear projections, each processed independently through a self-attention mechanism. The outputs of each attention head are then combined to obtain the final output of the multi-headed attention layer. Formally, taking the input modality with respect to a set of other modalities (), the multi-headed attention layer for OvO attention is defined as follows:
Here, is the number of attention heads, is a learnable weight matrix for the -th attention head, is a learnable weight matrix that projects the concatenated outputs of the attention heads back to the original dimension, and OvO Attention is defined in Equation 4.
3.3 Model Complexity
This section highlights the complexities of the two commonly used paradigms: early fusion followed by self-attention and pairwise cross-attention, as well as our One-Versus-Others (OvO) attention. Table 1 summarizes the complexity per layer. Let represent the number of modalities, be the feature-length of each modality (assuming equal), and be the representation dimension of the respective weight matrices. As established in (Vaswani et al., 2017), self-attention has complexity of . In the multimodal case, self-attention concatenates modalities before attention, leading to a sequence length of , influencing the quadratic term. Thus, the complexity of self-attention is . Cross-attention computes attention over all pairwise permutations of modalities: . Thus, the number of operations required by cross-attention is . When focusing on the fastest-growing terms in big notation, the final complexity per layer simplifies to . One-Versus-Others (OvO) Attention requires one attention calculation per modality, making it linear with respect to . Thus, the complexity per layer for OvO is . Appendix Section A provides step-by-step details for the complexity calculations.
| Model | Complexity Per Layer |
| Self-Attention | |
| Cross-Attention | |
| One-Versus-Others (OvO) Attention |
3.4 Illustration through simulation
To illustrate the linearity of OvO compared to the other integration paradigms, we simulated 20 artificial modalities. We consider two classes: (1) 20 random feature values that sum up to 1.0, and, (2) 20 random feature values that are each less than 0.15. These classes were created such that the correct label can only be inferred after inspecting all features. For example, 0.14 is less than 0.15, but it could also be a value that adds to 1. For more details on the simulation dataset and how the threshold was chosen, see Appendix B. Each value was then vectorized by sampling randomly around the selected number, such that each modality is a vector of size 20 rather than a single number, leading to a combined total of 400 features. Overall, the dataset contains 2,000 samples (1,000 for each class). Our constructed simulation dataset tests the scaling capabilities of our method to an extent that real-world datasets do not usually reach.
Using 2, 5, 10, 15, and 20 simulated modalities, we examine the computation cost across the three integration methods. Most notably, while self-attention and cross-attention grow quadratically with respect to the number of modalities, , (), our method scales linearly (), as shown in Figure 2.
4 Experiments
We used four diverse clinical datasets to examine our method against three standard integration techniques: concatenation with no attention (baseline), early fusion with self-attention, and pairwise cross-attention. These clinical tasks feature a range of rich modalities that, despite their high integration costs, remain essential to solve. In Section 5.1, we provide results on the Amazon Reviews dataset that show our method is applicable to fields outside the healthcare domain.
4.1 Dataset descriptions
4.1.1 The Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) dataset
The Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Petersen et al., 2010) database provides neuroimaging data, cognitive test scores, biomarker profiles, and genetic information for Alzheimer’s disease (AD), mild cognitive impairment (MCI), and normal patients. We use the processed data from the Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) challenge (Marinescu et al., 2019). We utilize six modalities that have the least missing information per patient: cognitive tests - neuropsychological tests administered by a clinical expert; MRI ROIs (generated from Freesurfer) - measures of brain structural integrity; FDG PET ROI averages - measure cell metabolism, where cells affected by AD show reduced metabolism; AV45 PET ROI averages - measures amyloid-beta load in the brain; demographic information (e.g., age, gender, education); and CSF biomarkers - amyloid and tau levels in the cerebrospinal fluid. The preprocessing provided by TADPOLE turned every modality into a tabular form (including imaging). After removing patients with missing modalities, we had 767 MCI patients, 493 normal patients, and 143 AD patients. Thus, this is a three-class classification task with six modalities.
4.1.2 The Cancer Genome Atlas (TCGA) dataset
TCGA is a cancer genomics program that molecularly characterized over 20,000 primary cancers and matched normal samples spanning 33 cancer types. Accurate prediction of cancer type is an important task but has primarily focused on single-modality approaches (Edge and Compton, 2010; Soh et al., 2017; Divate et al., 2022; Mohammed et al., 2021). Sun et al. (Sun et al., 2018) used multiple modalities from TCGA but did so for survival prediction rather than cancer classification. We used patients diagnosed with lung (Network et al., 2014b), stomach (Network et al., 2014a), colon (Network et al., 2012), liver (Ally et al., 2017), or kidney (working group: Baylor College of Medicine Creighton Chad J. 1 2 Morgan Margaret 1 Gunaratne Preethi H. 1 3 Wheeler David A. 1 Gibbs Richard A. 1 et al., 2013) cancer to create our dataset consisting of 5 modalities – whole-slide images, clinical data (tabular), copy number variation or CNV (tabular), gene expression (tabular), and DNA Methylation (tabular). CNV, DNA Methylation, and gene expression all describe different genomic information; full descriptions can be found in Appendix D. Clinical data includes demographics, laboratory tests, and family relationships. Imaging includes pathology slide images of tissue sampled from the tumor. In total, we had 338 colon cancer patients, 329 kidney, 301 lung, 228 liver, and 226 stomach patients after filtering described in Appendix D. This is a five-class classification task with five modalities.
4.1.3 MIMIC-IV and CXR
MIMIC-IV (Johnson et al., 2023) covers 431K visits for 180K patients admitted to the ICU in the Beth Israel Deaconess Medical Center. MIMIC Chest X-ray (MIMIC-CXR) (Johnson et al., 2019) contains chest radiographs in DICOM format with free-text radiology reports. The dataset contains 377,110 images corresponding to 227,835 radiographic studies performed at Beth Israel Medical Center. We follow the pre-processing of MedFuse (Mohammed et al., 2021) to extract the clinical time-series data from MIMIC-IV along with the associated chest X-ray images in MIMIC-CXR. We further expand the number of modalities by adding a demographics table and discharge notes, resulting in four modalities. We also follow MedFuse in the construction of the phenotyping task. The goal of this multi-label classification task is to predict whether a set of 25 chronic, mixed, and acute care conditions are assigned to a patient in a given ICU stay. This is a 25-class multi-label task with four modalities.
4.1.4 eICU
The eICU collaborative database encompasses data derived from intensive care units (ICU) across the continental United States between 2014 and 2015 (Pollard et al., 2018). The database is structured with a collection of tables relating to patients in the program, linked through a patient unit stay ID. For our predictive task aimed at determining mortality, we focus on six tabular modalities represented by the patient, diagnosis, treatment, medication, lab, and apacheApsVar tables. The apacheApsVar contains numerical variables used to calculate the Acute Physiology Score (APS) for patients, which summarizes a patient’s severity of illness. Details on other modalities can be found in Appendix C. After isolating ICU stays that had complete data for all modalities, our dataset comprises 75,845 unique patients associated with 93,784 ICU stays, of which 86,012 were recorded as alive and 7,772 were recorded as dead. This is a 2-class classification task with six modalities.
4.2 Baselines
Our multimodal baselines include a conventional concatenation fusion with no attention, early fusion followed by self-attention, and pairwise cross-attention fusion. The architectures of all models are identical except for their integration stage. For example, since modality-specific encoders can produce different dimension sizes, we add a linear layer before integration to create the same input dimensions. While this step is not strictly necessary for concatenation, we still add the layer there so that no additional factors influence computation costs and performance. While there are many multimodal Transformers available for the vision-language domain, our focus is on examining the underlying fusion mechanism and creating a general integration paradigm for any application, especially ones outside of vision-language.
4.3 Implementation details
For the MIMIC dataset, we follow the established train, validation, and test split in (Hayat et al., 2022). Similarly, for the TADPOLE task, we use the provided data splits, but add a constraint that repeating patients cannot appear across data splits, as to avoid information leakage. In the other datasets, for consistency, we randomly sampled 80% of the data for the training set and 10% each for test and validation sets, as there was not an established and publicly accepted split. To evaluate our model against other integration techniques, we use the domain-accepted metrics for each task: For MIMIC and eICU, we use area under the receiver operating characteristic (AUROC) and area under Precision-Recall (PR) curve (AUPRC) as established in past works (Hayat et al., 2022; Sheikhalishahi et al., 2020); For TADPOLE we use the multi-class area under the receiver operating curve (mAUC) and the overall balanced classification accuracy (BCA), as established by the competition creators (Marinescu et al., 2019); For TCGA, we use accuracy and F1-score as measures of performance, used in (Mohammed et al., 2021). For all datasets, we used the number of floating-point operations (FLOPs) as the measure of runtime complexity. FLOPs were measured per sample and reported as the difference between concatenation, the simplest integration setting, and multimodal attention (FLOPs).
4.4 Hyperparameter Tuning
Our hyperparameter tuning scheme was consistent for each dataset and each model. For each experiment, we used the evaluation metrics on the validation set to determine the best hyperparameters. We tuned the learning rate (0.01 - 1 x , dividing by 10 for each interval), batch size (16, 32, 64, 128), epochs (200 epochs with early stopping if validation performance did not increase for 5 epochs), and number of attention heads for OvO, self-attention, and cross-attention models (1, 2, 4, 8, 16). For the neural network encoders, we tuned the number of linear layers ranging from 1 to 4. Similarly, for the convolutional neural network, we tuned the number of convolution layers ranging from 1 to 4. For compute times and GPU details used for hyperparameter tuning, see Appendix G. Lastly, we randomly picked 10 random seeds for every experiment - once the best hyperparameters were picked, ten models initialized with those seeds and parameters were run. Then, using the trained ten models, we evaluated on the test set and took the average of the 10 runs along with the standard deviation, which is reported in Section 5.
5 Results
Using four real-world clinical datasets, diverse in terms of the number of modalities, feature space, and classification tasks, we demonstrate that our method consistently and drastically reduces computational costs compared to early fusion and pairwise fusion while simultaneously maintaining or enhancing performance. This is demonstrated on four-modality, five-modality, and two six-modality datasets.
For the four-modality MIMIC task, we used a pre-trained ClinicalBERT model for the text modality. For all other modalities, we used the appropriate neural network architecture (i.e., CNN for images, LSTM for time series, and a multi-layer perception for all tabular data). We perform significance testing between OvO attention’s and the next best performing model, detailed in the Appendix H.
The results on MIMIC are presented in Table 2, clearly demonstrating the scalability and performance advantages of OvO attention. OvO’s 4,227,072 FLOPs notably reduce computational costs compared to cross-attention (52,723,712 FLOPs) and self-attention (67,633,152 FLOPs), achieving reductions by 91.98% and 93.75%, respectively, thus highlighting OvO’s superior efficiency. The unimodal results show that the textual modality is most valuable in phenotype prediction, and ClinicalBERT alone performs better than self-attention and cross-attention. This indicates that the added complexity and forced interactions are not necessarily conducive to result quality. However, OvO attention can extract information from the other modalities for a significant performance increase rather than a decrease (p-value <0.01, see Appendix H).
| Model | Modalities | FLOPs | AUROC | AUPRC |
| LSTM | TS | - | 58.8 ±0.6 | 28.5 ±0.4 |
| CNN | I | - | 56.9 ±0.3 | 26.7 ±0.2 |
| NN | D | - | 64.1 ±0.4 | 32.4 ±0.3 |
| ClinicalBERT | T | - | 79.3 ±0.4 | 58.7 ±0.3 |
| Concatenation | All | * | 82.7 ±0.6 | 65.1 ±1.8 |
| Cross-Attention | All | 52,723,712 | 78.2 ±2.1 | 54.1 ±2.7 |
| Self-Attention | All | 67,633,152 | 78.5 ±2.0 | 55.7 ±3.1 |
| OvO Attention | All | 4,227,072 | 83.6 ±1.1 | 66.2 ±2.6 |
For the five-modality cancer prediction task from TCGA, we show our results in Table 3. While here, cross-attention yielded slightly better performance than OvO, our model offers substantial scalability benefits and, across 10 random seeds, performed consistently, showing robustness and stability. Furthermore, the difference between cross-attention and OvO attention accuracy and F1-scores was not statistically significant (p-value >0.01, see Appendix H). More importantly, OvO’s 330,240 FLOPs significantly undercut cross (5,260,800 FLOPs) and self-attention (6,556,160 FLOPs), achieving reductions of 93.73% and 94.96%, respectively, highlighting OvO’s superior efficiency.
| Model | Modalities | FLOPs | Accuracy | F1-Score |
| CNN | I | - | 56.7 ±3.0 | 55.6 ±4.0 |
| NN | C | - | 60.8 ±0.5 | 57.2 ±0.3 |
| NN | CNV | - | 93.4 ±0.8 | 94.1 ±0.7 |
| NN | M | - | 97.0 ±0.4 | 97.4 ±0.4 |
| NN | GE | - | 97.6 ±1.2 | 97.7 ±1.0 |
| Concatenation | All | * | 97.8 ±1.5 | 97.9 ±1.3 |
| Self-Attention | All | 6,556,160 | 96.8 ±1.0 | 97.0 ±0.9 |
| Cross-Attention | All | 5,260,800 | 99.2 ±1.0 | 99.3 ±0.9 |
| OvO Attention | All | 330,240 | 98.3 ±0.1 | 98.4 ±0.1 |
For the six-modality Alzheimer’s detection task from TADPOLE, we show our results in Table 4. OvO’s 405,504 FLOPs significantly undercut cross-attention (8,921,088 FLOPs) and self-attention (9,633,792 FLOPs), achieving reductions of 95.45% and 95.79%, respectively, highlighting OvO’s remarkable efficiency. Similarly to the MIMIC results, the unimodal results show that the cognitive tests modality is most valuable in disease prediction, and performs on its own better than self-attention and cross-attention. However, OvO attention can extract information from the other modalities for a significant performance increase rather than a decrease (p-value <0.01).
| Model | Modalities | FLOPs | MAUC | BCA |
| NN | A | - | 63.5 ±3.1 | 56.4 ±3.8 |
| NN | CSF | - | 64.4 ±1.1 | 53.6 ±2.7 |
| NN | M | - | 67.0 ±1.3 | 57.2 ±1.0 |
| NN | F | - | 66.6 ±0.3 | 60.8 ±0.7 |
| NN | D | - | 74.6 ±0.9 | 62.0 ±0.6 |
| NN | CT | - | 97.8 ±0.2 | 88.6 ±0.7 |
| Concatenation | All | * | 97.7 ±0.8 | 91.9 ±1.9 |
| Cross-Attention | All | 8,921,088 | 97.1 ±0.6 | 90.7 ±1.7 |
| Self-Attention | All | 9,633,792 | 94.8 ±1.1 | 86.6 ±2.6 |
| OvO Attention | All | 405,504 | 98.3 ±0.4 | 93.0 ±1.4 |
Lastly, the results on the six-modality eICU mortality prediction task are shown in Table 5, demonstrating the scalability and performance advantages of OvO attention. OvO’s 6,340,608 FLOPs significantly undercut those of cross-attention (129,957,888 FLOPs) and self-attention (151,781,376 FLOPs), achieving reductions of approximately 95.12% and 95.82%, respectively, thereby highlighting OvO’s efficiency. Mirroring the trends observed in the MIMIC and TADPOLE datasets, we note a dominant unimodal modality, specifically Lab modality, in this experiment as well. While concatenating modalities does enhance performance, this improvement is not seen in self and cross-attention models. In contrast, OvO attention not only reflects these performance gains but does so significantly (p-value <0.01). We hypothesize that this is due to the overfitting of more complex integration frameworks of self and cross-attention on relatively smaller clinical datasets. OvO, in its simplicity akin to concatenation, manages to strike a balance by maintaining flexibility and capturing inter-modal interactions through its attention mechanism, thus offering an edge in performance without excessive complexity.
| Model | Modalities | FLOPs | AUROC | AUPRC |
| NN | DE | - | 50.2 ±0.6 | 91.8 ±0.2 |
| NN | M | - | 56.3 ±1.3 | 93.1 ±0.3 |
| NN | D | - | 58.2 ±2.1 | 93.3 ±0.4 |
| NN | T | - | 66.1 ±0.5 | 94.8 ±0.1 |
| NN | A | - | 77.6 ±0.2 | 97.0 ±0.1 |
| NN | L | - | 81.5 ±0.4 | 97.0 ±0.1 |
| Concatenation | All | * | 81.7 ±1.6 | 97.5 ±0.3 |
| Cross-Attention | All | 129,957,888 | 77.6 ±1.6 | 95.4 ±0.3 |
| Self-Attention | All | 151,781,376 | 80.2 ±2.0 | 96.8 ±0.4 |
| OvO Attention | All | 6,340,608 | 82.5 ±0.9 | 97.8 ±0.2 |
In summary, across diverse clinical datasets and modalities, OvO attention consistently outperforms traditional fusion techniques in both predictive performance and computational efficiency, underlining its robustness and adaptability in handling complex multimodal healthcare data.
5.1 Generalizability
While our work focused on the clinical domain, where OvO will be most impactful and relevant, our method is a domain-agnostic approach and can be highly effective in other multimodal scenarios. We use the Amazon reviews dataset (Ni et al., 2019), to demonstrate OvO’s performance on a non-clinical dataset. The Amazon Reviews dataset includes review images, review text, and product metadata (e.g., price, category), with the goal of review sentiment classification (positive or negative). See Appendix F for more pre-processing details. The Amazon Reviews results are shown in Table 6 and demonstrate both the scalability and performance advantages of OvO attention. OvO’s 523,520 FLOPs significantly undercut those of cross-attention (1,903,616 FLOPs) and self-attention (2,685,440 FLOPs), achieving reductions of 72.50% and 80.51%, respectively, thereby highlighting OvO’s efficiency on the integration. Even though the efficiency gains scale with the number of modalities (see Figure 2), it is impressive to see significant improvements can be achieved for a dataset with just three modalities. Since the textual modality is most valuable in sentiment prediction, the performance of BERT alone is higher than concatenation, self-attention, and cross-attention. This indicates that the noise from metadata and images interferes with model performance. However, OvO attention can extract information from the other two modalities for a significant performance increase rather than a decrease (p-value <0.01).
| Model | Modalities | FLOPs | Accuracy | F1-Score |
| NN | Tb | - | 57.6 ±0.7 | 57.5 ±0.8 |
| ResNet | I | - | 66.3 ±0.7 | 66.6 ±0.6 |
| BERT | T | - | 92.6 ±0.5 | 92.9 ±0.4 |
| Concatenation | All | * | 92.2 ±0.4 | 92.8 ±0.3 |
| Self-Attention | All | 2,685,440 | 92.4 ±0.4 | 92.4 ±0.4 |
| Cross-Attention | All | 1,903,616 | 91.6 ±0.7 | 92.2 ±0.6 |
| OvO Attention | All | 523,520 | 93.1 ±0.3 | 93.0 ±0.3 |
Overall, the adaptability and performance enhancement demonstrated by OvO across various applications, not limited to the medical field, highlight its potential as a universal scalable solution in multimodal learning. OvO’s domain-agnostic approach can significantly contribute to overcoming the computational bottlenecks commonly encountered in multimodal scenarios.
6 Discussion
Our proposed method, OvO, provides a way to overcome one of the major challenges associated with multimodal datasets - computational resource demand and cost. Thus enabling adoption in resource-constrained domains, such as clinical decision support.
The literature on multimodal models and, specifically, Multimodal Transformers provides different architectures, some of which make efforts towards scalability. However, such models focus on the vision-language domain, and are not applicable in all multimodal scenarios (e.g., clinical decision support). Thus, we chose common fusion paradigms (similar to (Nagrani et al., 2021)) that encapsulate most of the underlying attention mechanisms in the multimodal literature. We aim to modify the basic attention mechanism in such models to make multimodal fusion more scalable with an increasing number of modalities.
7 Conclusion
We present One-Versus-Others (OvO), a new scalable multimodal attention mechanism. The proposed formulation significantly reduces the computational complexity compared to the widely used early fusion through self-attention and cross-attention methods. Notably, OvO achieves, at minimum, a reduction of 91.98% in FLOPs when benchmarked against self and cross-attention methods across a range of clinical datasets encompassing four, five, and six modalities. We provide both a detailed theoretical complexity analysis and empirical evidence from a simulated experiment, illustrating that OvO’s computational demand scales linearly with the number of modalities, in contrast to the quadratic scaling observed in other methods. Furthermore, the versatility of OvO is evidenced through its application to a non-clinical dataset, where it consistently surpassed all comparative baselines. Overall, the results unequivocally establish that OvO not only significantly reduces computational expenses but also meets or exceeds the performance of existing state-of-the-art fusion methodologies.
8 Impact Statement
Our proposed work, OvO, provides a method to overcome one of the major challenges associated with multimodal datasets - computational resource demand and cost. The scalability of the proposed formula allows diverse multimodal experiments to run in a wider range of research environments despite economic and computational constraints. While our method of modality fusion is meant to reduce the negative environmental impact of attention-learning models, we note that the individual modality encoders used (e.g., ClinicalBERT) still have a negative impact by creating emissions.
Appendix A Computational Complexity Analysis for Multimodal Integration Schemes
In this section, we present the step-by-step details of the computational complexity analysis presented in Section 3.3. The analysis is done with respect to the size of the input modalities associated with the three paradigms used in our experimental setting: early fusion followed by self-attention, cross-modal attention, and One-Versus-Others (OvO) Attention.
A.1 Early Fusion
The early fusion approach involves first combining the modalities and then processing the concatenated sequence with the self-attention mechanism.
Step 1: Concatenation of Modalities.
Let be the number of modalities and be the feature-length of each modality.
The complexity for this operation is linear:
Step 2: Compute Queries, Keys, and Values.
The self-attention mechanism derives queries (Q), keys (K), and values (V) for the concatenated sequence (length ) using linear transformations with representation dimension, . The complexity of each transformation operation is:
Step 3: Compute Attention Scores.
Attention scores are computed by taking the dot product of queries and keys. The self-attention mechanism has quadratic complexity with respect to the sequence length and linear complexity with respect to the representation dimension [Vaswani et al., 2017]. Thus, given the concatenated sequence’s length of and the dimension of the keys and queries , the complexity of this step is:
Step 4: Calculate the Weighted Sum for Outputs.
For each of the positions in the concatenated sequence, we compute the softmax of the attention scores to produce the attention weights. These weights are then multiplied with their corresponding -dimensional values to compute the weighted sum, which becomes the output. The computational complexity of these operations is:
When combining all steps, the dominating terms in the computational complexity stem from the attention scores’ computation and the weighted sum, culminating in an overall complexity of:
A.2 Cross-modal Attention
For cross-modal attention, each modality attends to every other modality.
Step 1: Compute Queries, Keys, and Values for Inter-Modal Attention.
From a given modality, compute a query (Q), and from the remaining modalities, compute keys (K) and values (V). Keys, queries, and values are obtained using linear transformations with representation dimension . The complexity of each transformation operation is:
Considering all modalities:
The term comes from the number of pairwise permutations of , given by .
Step 2: Calculate Attention Scores for Inter-Modal Attention.
The queries and keys from different modalities are used to compute attention scores, which represent how much one modality should attend to another.
Considering all modalities:
Step 3: Calculate the Weighted Sum for Outputs.
For every modality interaction, calculate the softmax of the attention scores to obtain the attention weights. These weights are then used in conjunction with the values vector to derive the weighted sum for the output:
Considering all modalities:
When evaluating all steps together, the dominating factors in computational complexity arise from the computation of attention scores and the weighted sum. Thus, the collective complexity for cross-modal attention, where each modality attends to every other, equates to:
For the complexity of cross-modal attention, the dominant term is . The term effectively becomes a constant factor in relation to . As tends toward larger values, the difference between and diminishes. This is a consequence of the principles of big notation, which focuses on the fastest-growing term in the equation while dismissing constant factors and lower-order terms. As a result, for asymptotic analysis, the complexity
can be simplified to:
A.3 One-Versus-Others (OvO) Attention Complexity
Step 1: Averaging of ”Other” Modalities.
Let be the number of modalities and be the feature-length of each modality. For each modality , averaging over the other modalities results in a complexity of:
Given that this needs to be computed for all modalities:
Step 2: Calculate Attention Scores with Shared Weight Matrix W.
The modality vector and the average of ”other” modalities, , are used to compute attention scores, which represent how much one modality should attend to the others. Multiplication with the weight matrix (with representation dimension ) and the dot product with the summed modalities lead to:
Considering this operation for all modalities:
Step 3: Calculate the Weighted Sum for Outputs.
For every modality interaction, calculate the softmax of the attention scores to obtain the attention weights. These weights are then used in conjunction with the vector (analogous the values (V) vector) to derive the weighted sum for the output:
Considering all modalities:
When evaluating all steps together, the dominating factors in computational complexity arise from the computation of attention scores. Thus, the collective complexity for cross-modal attention, where each modality attends to every other, equates to:
In summary, One-Versus-Others (OvO) Attention exhibits a computational complexity that grows linearly with respect to the number of modalities (). In contrast, both early fusion through self-attention and cross-attention approaches demonstrate quadratic growth with respect to the number of modalities (). This makes OvO a more scalable option for multimodal integration.
Appendix B Simulation Dataset Details
We consider two classes: (1) 20 random feature values that sum up to 1.0, and, (2) 20 random feature values that are each less than 0.15. The threshold was chosen at 0.15 because if 0.10 was the threshold, the mean of the 20 values would be 0.05, and thus, the sum would also be very close to 1, on average. This would render the task too difficult, and there would not be a significant difference between the samples across the two labels. Setting the threshold to 0.2 would render the task too easy, as on average, the numbers are consistently greater in the second class and the classes could be differentiated using only one modality. Thus, we chose 0.15 as the threshold.
Appendix C eICU Modality Descriptions and Detailed Pre-processing
The patient table contains patient demographic, admission, and discharge details. We use this table to determine mortality status. The diagnosis table contains active diagnoses given to each patient. We extract diagnosis features by one-hot encoding conditions specified in the ‘diagnosisstring’ column. The treatment table contains active treatments prescribed for each patient. We extract treatment features by one-hot encoding treatment types specified in the ‘treatmentstring’ column. The medication contains active medication orders for patients and when they were ordered. We extract medication features by one-hot encoding drugs specified in the ‘drugname’ column. The lab table contains lab type and corresponding results for various lab measurements collected for each patient. We extract lab features by summing lab measurements of commonly recorded lab types for each patient stay.
Appendix D TCGA Modality Descriptions and Detailed Pre-processing
CNV defines the varying number of repeats of genetic fragments found in a human genome. The number of repeats of specific genetic fragments influences gene expression levels and has been associated with the progression of different cancers [Samulin Erdem et al., 2017]. Any genomic regions missing CNV values or only having one unique value across all cases were removed. DNA methylation represents the amount of condensation of genetic regions due to the chemical alteration imposed by methyl groups. This condensation generally represses gene activity near the genetic region. Any genomic regions with missing values were removed. Clinical data includes information such as the patient’s diagnosis, demographics, laboratory tests, and family relationships. Categorical features were isolated and a coefficient of variation test was run to determine highly variable features. Features with a coefficient of variation higher than 70 were kept for analysis, along with the target variable. These features were converted into numerical format using one-hot-encoding. Gene expression data is collected through RNA-sequencing. Levels of gene expression are recorded by detecting the amounts of transcripts found for each gene. These levels can be used to determine the molecular mechanisms underlying cancer. Transcriptomic data was filtered to only include protein-coding genes and measured in fragments per kilobase of exon per million mapped fragments (FPKM). Imaging - TCGA collects pathology slide images of tissues sampled from the tumor. This modality provides visual information about the malignant region and can help with diagnosis and treatment planning. The image data was filtered only to include DX images, which result from a single X-Ray exposure, rotated to landscape view, then cropped to the median aspect ratio of 1.3565. We filtered for patients that had all five modalities, and we also only chose the patients that were still alive, to create a more balanced number of patients between cancer types (338 colon cancer patients, 329 kidney, 301 lung, 228 liver, and 226 stomach patients, after the filtering). The task we created is to classify each patient’s cancer type. For all modalities, features with missing values were dropped. For CNV, DNA Methylation, and gene expression data, feature reduction was performed using a random forest classifier, only on training data, ensuring the test was not seen by the random forest. Using the validation set, we determined the best number of estimators (out of 50, 100, 150).
Appendix E Hyperparameter Tuning
For each experiment, we used the validation accuracy to determine the best hyperparameters. We tuned the learning rate (0.01 - ), batch size (16, 32, 64, 128), epochs (200 epochs with early stopping if validation accuracy did not increase for 5 epochs), and number of attention heads for the OvO and pairwise cross-modal attention models (1, 2, 4, 8, 16). For the neural network encoders, we tuned the number of linear layers ranging from 1 to 4. Similarly, for the convolutional neural network, we tuned the number of convolution layers ranging from 1 to 4.
Appendix F Amazon Reviews Dataset Description and Detailed Pre-processing
The Amazon customer review dataset [Ni et al., 2019] aims to understand consumer opinions and experiences with products sold on Amazon. We sampled from the Electronics category, as it was one of the largest and most commonly used [Li et al., 2021, Haque et al., 2018, Pandey and Joshi, 2022]. We used three modalities from the dataset – review images, text, and the metadata (price, brand name, product category, etc.) associated with each review. The task is to classify the binary sentiment of each review - positive (4-5 stars) or negative (1-2 stars), with a total of around 20,000 samples. While sentiment classification with the Amazon reviews data is well-studied unimodally through text [Haque et al., 2018, Li et al., 2021, 2023, Pandey and Joshi, 2022], we offer a unique multimodal approach to the task. Existing multimodal approaches on Amazon reviews focus on review helpfulness [Han et al., 2022, Liu et al., 2021], or product similarity [Das et al., 2022] rather than sentiment classification. Furthermore, other papers use only text and images or only images and metadata (as text); an image-text-tabular multimodal task has yet to be explored. Intuitively, images and metadata can both be useful for review sentiment prediction - if the image highlights a broken item with the metadata showing a high price or trusted brand, consumer dissatisfaction may be implied.
Appendix G Compute Resources
For each experiment, we use one NVIDIA GeForce RTX 3090 GPU. For the Hateful Memes task, single-modality models ran for roughly 40 minutes, and multi-modal models ran for roughly 55 minutes on average. For the Amazon reviews task, the single modality pre-trained models ran for roughly 50 minutes, the single modality neural network ran for a minute, and the multi-modal models ran for approximately an hour on average. For the TCGA task, single-modality models ran for 5 minutes, while multi-modal models ran for roughly 15 minutes on average. In the simulation dataset, the maximum modalities was 20 which took our model, OvO, roughly 2 minutes to run, while the cross-modal attention baseline took about 20 minutes to run on average.
Appendix H Significance Testing
We use a t-test to determine if there is a significant difference the performance metrics (F1-score, accuracy, AUROC, AUPRC, MAUC, BCA) means between OvO attention and the next best-performing multimodal model. Our sample size is 10 from each group, as we initialized the models with 10 random seeds. For the MIMIC IV and CXR dataset, we compare against self-attention as it performed the second best after OvO. Using an = 0.01, we have evidence to reject the null hypothesis and conclude that there is a statistically significant difference in means between single-attention and OvO attention. The p-value for the AUROC scores is 0.00363 and the p-value for AUPRC is 0.000948. For the TCGA dataset, we do not have evidence to reject the null hypothesis and cannot say that the accuracy and F1-score means were different between OvO and cross-attention since the p-values were greater than = 0.01 (p-value of 0.04 for accuracy means, and p-value of 0.02 for F1-score means). This demonstrates that although cross-attention performed slightly better than OvO, it was not statistically significant. For the TADPOLE challenge, we compare against cross-attention as it performed the second best after OvO. We get a p-value for MAUC scores of and a p-value of for BCA. Thus, we demonstrate a statistically significant difference in accuracy and F1-score means between self-attention and OvO attention. Lastly, for the eICU dataset, we compare against cross-attention as it performed the second best after OvO. We get a p-value for AUROC scores of and a p-value of for AUPRC. Thus, we demonstrate a statistically significant difference in accuracy and F1-score means between self-attention and OvO attention.

### v4

@@
-One-Versus-Others Attention: Scalable Multimodal Integration for Clinical Data
+One-Versus-Others Attention: Scalable Multimodal Integration for Biomedical Data
 Abstract
-Multimodal models have become increasingly important as they surpass single-modality approaches on diverse tasks ranging from question-answering to autonomous driving.
+Multimodal models have become increasingly important as they surpass single-modality approaches on diverse tasks ranging from question-answering to disease diagnosis.
 Despite the importance of multimodal learning, existing efforts focus on vision-language applications, where the number of modalities rarely exceeds four (images, text, audio, video).
-However, data in other domains, such as healthcare, may include many more modalities like X-rays, PET scans, MRIs, genetic screening, genomic data, and clinical notes, creating a need for both efficient and accurate data integration.
+However, data in healthcare domain, may include many more modalities like X-rays, PET scans, MRIs, genetic screening, genomic data, and clinical notes, creating a need for both efficient and accurate data integration.
 Many state-of-the-art multimodal models rely on cross-attention or self-attention for effective data integration, which do not scale well for applications with more than two modalities.
@@
 To address this, we propose a new attention mechanism, One-Versus-Others (OvO) attention, that scales linearly with the number of modalities, thus offering a significant reduction in computational complexity compared to existing multimodal attention methods.
-Using four clinical datasets with diverse modalities, we show that our method decreases computation costs while maintaining or increasing performance compared to popular integration techniques.
-Across all clinical datasets, OvO reduced the number of required floating point operations (FLOPs) by at least 91.98%, demonstrating its significant impact on efficiency and enabling wider adaptation.
-111Code is available at https://anonymous.4open.science/r/OvO-13BD/.
+Using three clinical datasets with multiple diverse modalities, we show that our method decreases computation costs while maintaining or increasing performance compared to popular integration techniques.
+Across all clinical datasets, OvO reduced the number of required floating point operations (FLOPs) by at least 91.98%, demonstrating its significant impact on efficiency and enabling multi-modal predictions in healthcare.111Code and Appendix are available at https://github.com/rsinghlab/OvO
+keywords:
+Multimodal learning; deep learning; attention mechanism; clinical decision support.© 2024 The Authors.
+Open Access chapter published by World Scientific Publishing Company and distributed under the terms of the Creative Commons Attribution Non-Commercial (CC BY-NC) 4.0 License.
 1 Introduction
 Multimodal learning has emerged as a promising approach, which enables joint learning from multiple modalities of data (e.g., text and images).
-Combining different modalities allows for a more comprehensive and accurate understanding of tasks such as clinical decision support (Ming et al., 2022; Golovanevsky et al., 2022; Hayat et al., 2022), image and video captioning (Yu et al., 2019; Seo et al., 2022), audio-visual speech recognition (Sterpu et al., 2018), and sentiment analysis (Poria et al., 2018).
+Combining different modalities allows for a more comprehensive and accurate understanding of tasks such as clinical decision support [18, 5, 6], image and video captioning [33, 26], audio-visual speech recognition [30], and sentiment analysis [23].
 Multimodal learning has been explored through various methods in machine learning and deep learning.
@@
 However, these fusion paradigms lack a key component - capturing explicit interaction between modalities.
-For example, in detecting Alzheimer’s Disease, genetic features help reinforce and ground the clinical information and thus lead to more robust decision-making (Golovanevsky et al., 2022).
+For example, in detecting Alzheimer’s Disease, genetic features help reinforce and ground the clinical information and thus lead to more robust decision-making [5].
 Such relevant interactions can be captured through the attention mechanism.
-Popular multimodal models, such as LXMERT (Tan and Bansal, 2019) and ViLBERT (Lu et al., 2019), use a fusion method that captures interactions between modalities using cross-attention.
-On the other hand, models such as VisualBERT (Li et al., 2019) and VL-BERT (Su et al., 2019) use early fusion, where vision and language inputs are concatenated early to learn multimodal through self-attention.
+Popular multimodal models, such as LXMERT [31] and BLIP [12], use a fusion method that captures interactions between modalities using cross-attention.
+On the other hand, models such as VisualBERT [13] and LLaVA [14] use early fusion, where vision and language inputs are concatenated early to learn multimodal through self-attention.
+The clinical domain embraced these approaches, with multimodal models like MedFuseNet [27] and ARMOUR [15] using cross-attention for medical vision question answering and mortality prediction.
+In parallel, models such as BioViL-T [2] and MMBERT [10] employ early fusion through self-attention for disease prediction and report generation.
 However, both self-attention and cross-attention grow quadratically in computational burden with the number of modalities, posing a scalability challenge.
@@
 Figure 1 sketches a four-modality example to demonstrate the difference between our approach (scales linearly) and self-attention/cross-attention (scales quadratically).
-OvO is a general attention scheme that can be integrated into existing multimodal architectures instead of cross-attention or self-attention.
-We first, present a complexity analysis and demonstrate its validity through a simulated dataset.
+OvO is a general attention scheme that can be integrated into existing clinical multimodal architectures instead of cross-attention or self-attention.
+We first, present a complexity analysis and validate it through a simulated dataset.
 Our simulation results show scalability gains in an extreme multimodal setting (20 modalities).
-Next, we use four diverse clinical datasets that vary in modalities, encoder types (pre-trained and not), number of samples, and tasks to show our model’s improved scalability in different clinical multimodal settings.
+Next, we use three diverse clinical datasets that vary in modalities, encoder types (pre-trained and not), number of samples, and tasks to show our model’s improved scalability in different clinical multimodal settings.
 Our results demonstrate that our method dramatically decreases computation costs (offering at minimum a 91.98% reduction in computations), compared to self-attention and cross-attention while maintaining or even exceeding performance.
-Finally, we demonstrate the generalizability of our method on an additional non-clinical dataset, further underscoring OvO’s broad applicability.
-Overall, OvO is a novel attention scheme for multimodal integration that is domain-agnostic and scales linearly to the number of modalities.
-Importantly, it enables the practical application of deep learning models in healthcare, where computational efficiency and accuracy are vital for deployment.
+Overall, OvO is a novel attention scheme for multimodal integration that scales linearly to the number of modalities, enabling the practical application of deep learning models in healthcare, where computational efficiency and accuracy are vital for deployment.
 2 Related work
 Multimodal attention-based models are increasingly pivotal in clinical decision support systems, paralleling their widespread use in vision-language applications.
-In the medical domain, these models have shown remarkable utility in diverse scenarios, such as cancer classification (LI et al., 2021), biomarker discovery (Braman et al., 2021; Ilyin et al., 2004), prognosis prediction (Schulz et al., 2021; Silva and Rohr, 2020), and more.
+In the medical domain, these models have shown remarkable utility in diverse scenarios, such as cancer classification [11], biomarker discovery [3, 7], prognosis prediction [25, 29], and more.
 These applications highlight the versatility and potential of multimodal learning in handling complex and rich medical data.
@@
 Commonly, the representations from each modality in the multimodal models are passed through one of two paradigms - early fusion followed by self-attention or fusion through cross-attention.
-The early fusion group (e.g., Transformer-based models like Uniter (Chen et al., 2020), VisualBERT (Li et al., 2019), Vl-BERT (Su et al., 2019)) concatenates the visual embeddings and the textual embeddings as a single input, before passing the inputs through attention (see Figure 1 (a)).
+The early fusion group (e.g., Transformer-based models like UNITER [4], VisualBERT [13], LLaVA [14], BioViL-T [2], MedViLL [20], etc.) concatenates the visual embeddings and the textual embeddings as a single input, before passing the inputs through attention (see Figure 1 (a)).
 Given modalities and , queries (), keys (), and values () are computed from their concatenated sequence (e.g., ).
 The final output from a standard Transformer block is denoted by , Equation 1 shows the early fusion paradigm.
-The cross-attention scheme (used in Transformer-based models like ViLBERT (Lu et al., 2019), LXMERT (Tan and Bansal, 2019), ActBERT (Zhu and Yang, 2020), MulT (Tsai et al., 2019), etc.) inputs each modality into its own Transformer, the outputs of which are fed to a cross-modal Transformer (see Figure 1 (b)).
+The cross-attention scheme (used in Transformer-based models like ViLBERT [16], LXMERT [31], MedFuseNet [27], MADDi [5], etc.) inputs each modality into its own Transformer, the outputs of which are fed to a cross-modal Transformer (see Figure 1 (b)).
 For such models, the cross-modal interactions are captured in a pairwise manner through cross-attention, where queries (), keys (), and values () are computed from the modality inputs ( and ), and then the keys and values from each modality are fed to the multi-headed attention block of the other modality.
 The output, , is shown in Equation 2.
-While the early fusion and cross-attention paradigms could be extended to three modalities, seen in TriBERT (Rahman et al., 2021) and VATT (Akbari et al., 2021), these models face scalability challenges for more than three modalities.
+While the early fusion and cross-attention paradigms could be extended to three modalities, seen in TriBERT [24] and VATT [1], these models face scalability challenges for more than three modalities.
 Cross-attention methods can leverage joint representations formed from cross-attention but do not scale well to larger numbers of modalities as they are computed in a pairwise fashion.
@@
 Early fusion involves the concatenation of modalities before the Transformer layer, which similarly does not scale well with the number of modalities.
-Self-attention is quadratic with respect to sequence length (Vaswani et al., 2017), and since early fusion methods concatenate inputs before attention, the computational complexity will increase quadratically as the number of modalities increases (see Section 3.3).
+Self-attention is quadratic with respect to sequence length [32], and since early fusion methods concatenate inputs before attention, the computational complexity will increase quadratically as the number of modalities increases (see Section 3.3).
 Furthermore, concatenation is not order invariant, making the ordering of modalities an important consideration, potentially requiring similar bi-directional computations as cross-attention.
@@
 3.1 One-Versus-Others (OvO) attention
-We propose a new attention mechanism, One-Versus-Others (OvO) Attention, which grows linearly with number of modalities rather than quadratically, as would be the case for cross-attention or self-attention (see Section A).
+We propose a new attention mechanism, One-Versus-Others (OvO) Attention, which grows linearly with the number of modalities rather than quadratically, as would be the case for cross-attention or self-attention (see Section 3.3).
 OvO computes attention between one modality at a time with respect to all other modalities.
-Given modality , which is an embedding obtained from a dedicated encoder, where is the number of modalities and , OvO takes in one modality and computes the dot product against all the other modalities with a weight matrix .
-is a learnable parameter that is shared across all modalities (see Figure 1 (c)) and can learn interactions throughout training.
-The similarity score function, representing the degree of alignment between the chosen modality and others, calculated for modality with respect to a set of other modalities () is shown in Equation 3.
+Given modality , which is an embedding obtained from a dedicated encoder (e.g., CNN, ClinicalBERT, etc.) and where is the number of modalities, OvO takes in one modality and computes the dot product against all the other modalities with a weight matrix .
+is a learnable parameter that can help scale the importance of each attention calculation (see Figure 1 (c)) and can learn interactions throughout training.
+The modality embeddings in OvO attention function like queries, while the weighted sum of the other modalities behaves like keys and values, akin to the dot-product attention mechanisms, with the naming choice of adapted for multimodal applications.
+This dot product The similarity score function, representing the degree of alignment between the chosen modality and others, calculated for modality with respect to a set of other modalities () is shown in Equation 3.
+This produces a vector of scores which measure the relevance of with respect to the other modalities.
 The context vector in OvO for modality , which is a combined representation of information from the other modalities, is shown in Equation 4:
-We chose to sum over the “other” modalities instead of concatenation for two reasons: (1) the concatenation vector will continue to increase in length with the number of modalities, which will result in a less scalable framework; (2) concatenation is not invariant to the order of modalities, which could affect the model prediction, whereas a sum provides position invariance.
-Note that for modalities (, ), the similarity score function simplifies to that of general attention (Luong et al., 2015):
+In Equation 4, the softmax is applied across the input dimension of the embeddings, normalizing across the attention scores.
+The result is then multiplied by the original modality embeddings to compute the final output.
+We chose to sum over the “other” modalities instead of concatenation because: (1) the concatenation vector will continue to increase in length with the number of modalities, which will result in a less scalable framework; (2) concatenation is not invariant to the order of modalities, which could affect the model prediction, whereas a sum provides position invariance.
+Unlike cross-attention and self-attention, OvO provides a more interpretable mechanism for analyzing interactions between multiple modalities.
+In cross-attention, interactions are captured in isolated pairs (e.g., compared to or compared to ), limiting the ability to see how one modality integrates information from all others.
+Similarly, self-attention condenses modality interactions into a single operation, which can obscure explicit cross-modal interactions and make it difficult to disentangle their contributions.
+OvO, however, creates one attention matrix per modality, with each modality interacting with all others through the learnable weight matrix , which ensures a flexible and adaptive scaling mechanism.
+In future work, will be studied to better understand the relative importance of each modality’s contributions to the final prediction.
 3.2 Multi-headed OvO Attention
@@
 Formally, taking the input modality with respect to a set of other modalities (), the multi-headed attention layer for OvO attention is defined as follows:
-Here, is the number of attention heads, is a learnable weight matrix for the -th attention head, is a learnable weight matrix that projects the concatenated outputs of the attention heads back to the original dimension, and OvO Attention is defined in Equation 4.
+Here, is the number of attention heads, is a learnable weight matrix for the -th attention head, is a learnable weight matrix that projects the concatenated outputs of the attention heads back to the original dimension, and Equation 4 defines OvO Attention.
 3.3 Model Complexity
 This section highlights the complexities of the two commonly used paradigms: early fusion followed by self-attention and pairwise cross-attention, as well as our One-Versus-Others (OvO) attention.
-Table 1 summarizes the complexity per layer.
+Table 3.3 summarizes the complexity per layer.
 Let represent the number of modalities, be the feature-length of each modality (assuming equal), and be the representation dimension of the respective weight matrices.
-As established in (Vaswani et al., 2017), self-attention has complexity of .
+As established in [32], self-attention has complexity of .
 In the multimodal case, self-attention concatenates modalities before attention, leading to a sequence length of , influencing the quadratic term.
@@
 Thus, the complexity per layer for OvO is .
-Appendix Section A provides step-by-step details for the complexity calculations.
-| Model | Complexity Per Layer |
-| Self-Attention | |
-| Cross-Attention | |
-| One-Versus-Others (OvO) Attention |
+Appendix Section 1 provides step-by-step details for the complexity calculations.
+[…]
 3.4 Illustration through simulation
@@
 For example, 0.14 is less than 0.15, but it could also be a value that adds to 1.
-For more details on the simulation dataset and how the threshold was chosen, see Appendix B.
+For more details on the simulation dataset and how the threshold was chosen, see Appendix Section 2.
 Each value was then vectorized by sampling randomly around the selected number, such that each modality is a vector of size 20 rather than a single number, leading to a combined total of 400 features.
@@
 Our constructed simulation dataset tests the scaling capabilities of our method to an extent that real-world datasets do not usually reach.
-Using 2, 5, 10, 15, and 20 simulated modalities, we examine the computation cost across the three integration methods.
+We examine the computation cost across the three integration methods using 2, 5, 10, 15, and 20 simulated modalities.
 Most notably, while self-attention and cross-attention grow quadratically with respect to the number of modalities, , (), our method scales linearly (), as shown in Figure 2.
 4 Experiments
-We used four diverse clinical datasets to examine our method against three standard integration techniques: concatenation with no attention (baseline), early fusion with self-attention, and pairwise cross-attention.
+We used three diverse clinical datasets to examine our method against three standard integration techniques: concatenation with no attention (baseline), early fusion with self-attention, and pairwise cross-attention.
 These clinical tasks feature a range of rich modalities that, despite their high integration costs, remain essential to solve.
-In Section 5.1, we provide results on the Amazon Reviews dataset that show our method is applicable to fields outside the healthcare domain.
 4.1 Dataset descriptions
-4.1.1 The Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) dataset
-The Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Petersen et al., 2010) database provides neuroimaging data, cognitive test scores, biomarker profiles, and genetic information for Alzheimer’s disease (AD), mild cognitive impairment (MCI), and normal patients.
-We use the processed data from the Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) challenge (Marinescu et al., 2019).
+4.1.1 MIMIC-IV and CXR data
+MIMIC-IV [9] covers 431K visits for 180K patients admitted to the ICU in the Beth Israel Deaconess Medical Center.
+MIMIC Chest X-ray (MIMIC-CXR) [8] contains chest radiographs in DICOM format with free-text radiology reports.
+The dataset contains 377,110 images corresponding to 227,835 radiographic studies performed at Beth Israel Medical Center.
+We follow the pre-processing of MedFuse [19] to extract the clinical time-series data from MIMIC-IV along with the associated chest X-ray images in MIMIC-CXR.
+We further expand the number of modalities by adding a demographics table and discharge notes, resulting in four modalities.
+We also follow MedFuse in the construction of the phenotyping task.
+The goal of this multi-label classification task is to predict whether a set of 25 chronic, mixed, and acute care conditions are assigned to a patient in a given ICU stay.
+This is a 25-class multi-label task with four modalities.
+4.1.2 The Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) data
+The Alzheimer’s Disease Neuroimaging Initiative (ADNI) [21] database provides neuroimaging data, cognitive test scores, biomarker profiles, and genetic information for Alzheimer’s disease (AD), mild cognitive impairment (MCI), and normal patients.
+We use the processed data from the Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) challenge [17].
+We focus on a one-time diagnosis prediction task, using the most recent available data point for each patient across all modalities.
+This ensures that each data sample includes information from the same time point, which aligns with our goal of evaluating cross-modal integration rather than longitudinal progression.
 We utilize six modalities that have the least missing information per patient: cognitive tests - neuropsychological tests administered by a clinical expert; MRI ROIs (generated from Freesurfer) - measures of brain structural integrity; FDG PET ROI averages - measure cell metabolism, where cells affected by AD show reduced metabolism; AV45 PET ROI averages - measures amyloid-beta load in the brain; demographic information (e.g., age, gender, education); and CSF biomarkers - amyloid and tau levels in the cerebrospinal fluid.
@@
 Thus, this is a three-class classification task with six modalities.
-4.1.2 The Cancer Genome Atlas (TCGA) dataset
-TCGA is a cancer genomics program that molecularly characterized over 20,000 primary cancers and matched normal samples spanning 33 cancer types.
-Accurate prediction of cancer type is an important task but has primarily focused on single-modality approaches (Edge and Compton, 2010; Soh et al., 2017; Divate et al., 2022; Mohammed et al., 2021).
-Sun et al.
-(Sun et al., 2018) used multiple modalities from TCGA but did so for survival prediction rather than cancer classification.
-We used patients diagnosed with lung (Network et al., 2014b), stomach (Network et al., 2014a), colon (Network et al., 2012), liver (Ally et al., 2017), or kidney (working group: Baylor College of Medicine Creighton Chad J.
-1 2 Morgan Margaret 1 Gunaratne Preethi H.
-1 3 Wheeler David A.
-1 Gibbs Richard A.
-1 et al., 2013) cancer to create our dataset consisting of 5 modalities – whole-slide images, clinical data (tabular), copy number variation or CNV (tabular), gene expression (tabular), and DNA Methylation (tabular).
-CNV, DNA Methylation, and gene expression all describe different genomic information; full descriptions can be found in Appendix D.
-Clinical data includes demographics, laboratory tests, and family relationships.
-Imaging includes pathology slide images of tissue sampled from the tumor.
-In total, we had 338 colon cancer patients, 329 kidney, 301 lung, 228 liver, and 226 stomach patients after filtering described in Appendix D.
-This is a five-class classification task with five modalities.
-4.1.3 MIMIC-IV and CXR
-MIMIC-IV (Johnson et al., 2023) covers 431K visits for 180K patients admitted to the ICU in the Beth Israel Deaconess Medical Center.
-MIMIC Chest X-ray (MIMIC-CXR) (Johnson et al., 2019) contains chest radiographs in DICOM format with free-text radiology reports.
-The dataset contains 377,110 images corresponding to 227,835 radiographic studies performed at Beth Israel Medical Center.
-We follow the pre-processing of MedFuse (Mohammed et al., 2021) to extract the clinical time-series data from MIMIC-IV along with the associated chest X-ray images in MIMIC-CXR.
-We further expand the number of modalities by adding a demographics table and discharge notes, resulting in four modalities.
-We also follow MedFuse in the construction of the phenotyping task.
-The goal of this multi-label classification task is to predict whether a set of 25 chronic, mixed, and acute care conditions are assigned to a patient in a given ICU stay.
-This is a 25-class multi-label task with four modalities.
-4.1.4 eICU
-The eICU collaborative database encompasses data derived from intensive care units (ICU) across the continental United States between 2014 and 2015 (Pollard et al., 2018).
-The database is structured with a collection of tables relating to patients in the program, linked through a patient unit stay ID.
-For our predictive task aimed at determining mortality, we focus on six tabular modalities represented by the patient, diagnosis, treatment, medication, lab, and apacheApsVar tables.
-The apacheApsVar contains numerical variables used to calculate the Acute Physiology Score (APS) for patients, which summarizes a patient’s severity of illness.
-Details on other modalities can be found in Appendix C.
-After isolating ICU stays that had complete data for all modalities, our dataset comprises 75,845 unique patients associated with 93,784 ICU stays, of which 86,012 were recorded as alive and 7,772 were recorded as dead.
+4.1.3 eICU data
+The eICU collaborative database includes data from ICUs across the continental United States between 2014 and 2015 [22].
+It consists of tables linked through a patient unit stay ID.
+For our mortality prediction task, we focus on six tabular modalities: patient, diagnosis, treatment, medication, lab, and apacheApsVar tables.
+The apacheApsVar table contains numerical variables used to calculate the Acute Physiology Score (APS), an established method within the Acute Physiology Age Chronic Health Evaluation (APACHE) system for summarizing patient’s severity of illness on ICU admission and predicting outcomes.
+The patient table includes demographic, admission, and discharge details, and is used to determine mortality status.
+The diagnosis table lists active diagnoses for each patient, the treatment table includes active treatments, and the medication table contains active medication orders.
+We extract the features from these tables by one-hot encoding the relevant conditions, treatment types, and drug names, respectively.
+The lab table includes lab results, with features extracted by summing commonly recorded lab types.
+Our dataset includes 75,845 unique patients with 93,784 ICU stays, 86,012 recorded as alive and 7,772 as dead.
 This is a 2-class classification task with six modalities.
@@
 For example, since modality-specific encoders can produce different dimension sizes, we add a linear layer before integration to create the same input dimensions.
-While this step is not strictly necessary for concatenation, we still add the layer there so that no additional factors influence computation costs and performance.
+Although this step is not strictly necessary for concatenation, we still add the layer there so that no additional factors influence computation costs and performance.
 While there are many multimodal Transformers available for the vision-language domain, our focus is on examining the underlying fusion mechanism and creating a general integration paradigm for any application, especially ones outside of vision-language.
+In Appendix Section 5, we touch on the limitations of our experiments and future work that we did not cover.
 4.3 Implementation details
-For the MIMIC dataset, we follow the established train, validation, and test split in (Hayat et al., 2022).
-Similarly, for the TADPOLE task, we use the provided data splits, but add a constraint that repeating patients cannot appear across data splits, as to avoid information leakage.
-In the other datasets, for consistency, we randomly sampled 80% of the data for the training set and 10% each for test and validation sets, as there was not an established and publicly accepted split.
-To evaluate our model against other integration techniques, we use the domain-accepted metrics for each task: For MIMIC and eICU, we use area under the receiver operating characteristic (AUROC) and area under Precision-Recall (PR) curve (AUPRC) as established in past works (Hayat et al., 2022; Sheikhalishahi et al., 2020); For TADPOLE we use the multi-class area under the receiver operating curve (mAUC) and the overall balanced classification accuracy (BCA), as established by the competition creators (Marinescu et al., 2019); For TCGA, we use accuracy and F1-score as measures of performance, used in (Mohammed et al., 2021).
+For the MIMIC dataset, we follow the established train, validation, and test split in Hayat et al [6].
+Similarly, for the TADPOLE task, we use the provided data splits but add a constraint that repeating patients cannot appear across data splits to avoid information leakage.
+In the other datasets, for consistency, we randomly sampled 80% of the data for the training set and 10% each for test and validation sets, as there was not an established split.
+To evaluate our model against other integration techniques, we use the domain-accepted metrics for each task: For MIMIC and eICU, we use area under the receiver operating characteristic (AUROC) and area under Precision-Recall (PR) curve (AUPRC) as established in past works [6, 28]; For TADPOLE we use the multi-class area under the receiver operating curve (mAUC) and the overall balanced classification accuracy (BCA), as established by the competition creators [17].
 For all datasets, we used the number of floating-point operations (FLOPs) as the measure of runtime complexity.
@@
 Similarly, for the convolutional neural network, we tuned the number of convolution layers ranging from 1 to 4.
-For compute times and GPU details used for hyperparameter tuning, see Appendix G.
+For compute times and GPU details used for hyperparameter tuning, see Appendix Section 2.
 Lastly, we randomly picked 10 random seeds for every experiment - once the best hyperparameters were picked, ten models initialized with those seeds and parameters were run.
@@
 5 Results
-Using four real-world clinical datasets, diverse in terms of the number of modalities, feature space, and classification tasks, we demonstrate that our method consistently and drastically reduces computational costs compared to early fusion and pairwise fusion while simultaneously maintaining or enhancing performance.
-This is demonstrated on four-modality, five-modality, and two six-modality datasets.
-For the four-modality MIMIC task, we used a pre-trained ClinicalBERT model for the text modality.
+Using three real-world clinical datasets, diverse in terms of the number of modalities, feature space, and classification tasks, we demonstrate that our method consistently and drastically reduces computational costs compared to early fusion and pairwise fusion while simultaneously maintaining or enhancing performance.
+This is demonstrated on one four-modality dataset and two six-modality datasets.
+For the four-modality MIMIC task, we used pre-trained ClinicalBERT model for the text modality and fine-tuned it for the unimodal baseline and the multimodal task, separately, ensuring adaptation in each setting.
 For all other modalities, we used the appropriate neural network architecture (i.e., CNN for images, LSTM for time series, and a multi-layer perception for all tabular data).
-We perform significance testing between OvO attention’s and the next best performing model, detailed in the Appendix H.
+We perform significance testing between OvO attention and the next best-performing model, detailed in Appendix Section 3.
+[…]
+[…]
 The results on MIMIC are presented in Table 5, clearly demonstrating the scalability and performance advantages of OvO attention.
@@
 This indicates that the added complexity and forced interactions are not necessarily conducive to result quality.
-However, OvO attention can extract information from the other modalities for a significant performance increase rather than a decrease (p-value <0.01, see Appendix H).
-| Model | Modalities | FLOPs | AUROC | AUPRC |
-| LSTM | TS | - | 58.8 ±0.6 | 28.5 ±0.4 |
-| CNN | I | - | 56.9 ±0.3 | 26.7 ±0.2 |
-| NN | D | - | 64.1 ±0.4 | 32.4 ±0.3 |
-| ClinicalBERT | T | - | 79.3 ±0.4 | 58.7 ±0.3 |
-| Concatenation | All | * | 82.7 ±0.6 | 65.1 ±1.8 |
-| Cross-Attention | All | 52,723,712 | 78.2 ±2.1 | 54.1 ±2.7 |
-| Self-Attention | All | 67,633,152 | 78.5 ±2.0 | 55.7 ±3.1 |
-| OvO Attention | All | 4,227,072 | 83.6 ±1.1 | 66.2 ±2.6 |
-For the five-modality cancer prediction task from TCGA, we show our results in Table 3.
-While here, cross-attention yielded slightly better performance than OvO, our model offers substantial scalability benefits and, across 10 random seeds, performed consistently, showing robustness and stability.
-Furthermore, the difference between cross-attention and OvO attention accuracy and F1-scores was not statistically significant (p-value >0.01, see Appendix H).
-More importantly, OvO’s 330,240 FLOPs significantly undercut cross (5,260,800 FLOPs) and self-attention (6,556,160 FLOPs), achieving reductions of 93.73% and 94.96%, respectively, highlighting OvO’s superior efficiency.
-| Model | Modalities | FLOPs | Accuracy | F1-Score |
-| CNN | I | - | 56.7 ±3.0 | 55.6 ±4.0 |
-| NN | C | - | 60.8 ±0.5 | 57.2 ±0.3 |
-| NN | CNV | - | 93.4 ±0.8 | 94.1 ±0.7 |
-| NN | M | - | 97.0 ±0.4 | 97.4 ±0.4 |
-| NN | GE | - | 97.6 ±1.2 | 97.7 ±1.0 |
-| Concatenation | All | * | 97.8 ±1.5 | 97.9 ±1.3 |
-| Self-Attention | All | 6,556,160 | 96.8 ±1.0 | 97.0 ±0.9 |
-| Cross-Attention | All | 5,260,800 | 99.2 ±1.0 | 99.3 ±0.9 |
-| OvO Attention | All | 330,240 | 98.3 ±0.1 | 98.4 ±0.1 |
+However, OvO attention can extract information from the other modalities for a significant performance increase rather than a decrease (p-value <0.01, see Appendix Section 3).
 For the six-modality Alzheimer’s detection task from TADPOLE, we show our results in Table 5.
@@
 However, OvO attention can extract information from the other modalities for a significant performance increase rather than a decrease (p-value <0.01).
-| Model | Modalities | FLOPs | MAUC | BCA |
-| NN | A | - | 63.5 ±3.1 | 56.4 ±3.8 |
-| NN | CSF | - | 64.4 ±1.1 | 53.6 ±2.7 |
-| NN | M | - | 67.0 ±1.3 | 57.2 ±1.0 |
-| NN | F | - | 66.6 ±0.3 | 60.8 ±0.7 |
-| NN | D | - | 74.6 ±0.9 | 62.0 ±0.6 |
-| NN | CT | - | 97.8 ±0.2 | 88.6 ±0.7 |
-| Concatenation | All | * | 97.7 ±0.8 | 91.9 ±1.9 |
-| Cross-Attention | All | 8,921,088 | 97.1 ±0.6 | 90.7 ±1.7 |
-| Self-Attention | All | 9,633,792 | 94.8 ±1.1 | 86.6 ±2.6 |
-| OvO Attention | All | 405,504 | 98.3 ±0.4 | 93.0 ±1.4 |
 Lastly, the results on the six-modality eICU mortality prediction task are shown in Table 5, demonstrating the scalability and performance advantages of OvO attention.
+[…]
 OvO’s 6,340,608 FLOPs significantly undercut those of cross-attention (129,957,888 FLOPs) and self-attention (151,781,376 FLOPs), achieving reductions of approximately 95.12% and 95.82%, respectively, thereby highlighting OvO’s efficiency.
@@
 OvO, in its simplicity akin to concatenation, manages to strike a balance by maintaining flexibility and capturing inter-modal interactions through its attention mechanism, thus offering an edge in performance without excessive complexity.
-| Model | Modalities | FLOPs | AUROC | AUPRC |
-| NN | DE | - | 50.2 ±0.6 | 91.8 ±0.2 |
-| NN | M | - | 56.3 ±1.3 | 93.1 ±0.3 |
-| NN | D | - | 58.2 ±2.1 | 93.3 ±0.4 |
-| NN | T | - | 66.1 ±0.5 | 94.8 ±0.1 |
-| NN | A | - | 77.6 ±0.2 | 97.0 ±0.1 |
-| NN | L | - | 81.5 ±0.4 | 97.0 ±0.1 |
-| Concatenation | All | * | 81.7 ±1.6 | 97.5 ±0.3 |
-| Cross-Attention | All | 129,957,888 | 77.6 ±1.6 | 95.4 ±0.3 |
-| Self-Attention | All | 151,781,376 | 80.2 ±2.0 | 96.8 ±0.4 |
-| OvO Attention | All | 6,340,608 | 82.5 ±0.9 | 97.8 ±0.2 |
-In summary, across diverse clinical datasets and modalities, OvO attention consistently outperforms traditional fusion techniques in both predictive performance and computational efficiency, underlining its robustness and adaptability in handling complex multimodal healthcare data.
-5.1 Generalizability
-While our work focused on the clinical domain, where OvO will be most impactful and relevant, our method is a domain-agnostic approach and can be highly effective in other multimodal scenarios.
-We use the Amazon reviews dataset (Ni et al., 2019), to demonstrate OvO’s performance on a non-clinical dataset.
-The Amazon Reviews dataset includes review images, review text, and product metadata (e.g., price, category), with the goal of review sentiment classification (positive or negative).
-See Appendix F for more pre-processing details.
-The Amazon Reviews results are shown in Table 6 and demonstrate both the scalability and performance advantages of OvO attention.
-OvO’s 523,520 FLOPs significantly undercut those of cross-attention (1,903,616 FLOPs) and self-attention (2,685,440 FLOPs), achieving reductions of 72.50% and 80.51%, respectively, thereby highlighting OvO’s efficiency on the integration.
-Even though the efficiency gains scale with the number of modalities (see Figure 2), it is impressive to see significant improvements can be achieved for a dataset with just three modalities.
-Since the textual modality is most valuable in sentiment prediction, the performance of BERT alone is higher than concatenation, self-attention, and cross-attention.
-This indicates that the noise from metadata and images interferes with model performance.
-However, OvO attention can extract information from the other two modalities for a significant performance increase rather than a decrease (p-value <0.01).
-| Model | Modalities | FLOPs | Accuracy | F1-Score |
-| NN | Tb | - | 57.6 ±0.7 | 57.5 ±0.8 |
-| ResNet | I | - | 66.3 ±0.7 | 66.6 ±0.6 |
-| BERT | T | - | 92.6 ±0.5 | 92.9 ±0.4 |
-| Concatenation | All | * | 92.2 ±0.4 | 92.8 ±0.3 |
-| Self-Attention | All | 2,685,440 | 92.4 ±0.4 | 92.4 ±0.4 |
-| Cross-Attention | All | 1,903,616 | 91.6 ±0.7 | 92.2 ±0.6 |
-| OvO Attention | All | 523,520 | 93.1 ±0.3 | 93.0 ±0.3 |
-Overall, the adaptability and performance enhancement demonstrated by OvO across various applications, not limited to the medical field, highlight its potential as a universal scalable solution in multimodal learning.
-OvO’s domain-agnostic approach can significantly contribute to overcoming the computational bottlenecks commonly encountered in multimodal scenarios.
-6 Discussion
-Our proposed method, OvO, provides a way to overcome one of the major challenges associated with multimodal datasets - computational resource demand and cost.
-Thus enabling adoption in resource-constrained domains, such as clinical decision support.
-The literature on multimodal models and, specifically, Multimodal Transformers provides different architectures, some of which make efforts towards scalability.
-However, such models focus on the vision-language domain, and are not applicable in all multimodal scenarios (e.g., clinical decision support).
-Thus, we chose common fusion paradigms (similar to (Nagrani et al., 2021)) that encapsulate most of the underlying attention mechanisms in the multimodal literature.
-We aim to modify the basic attention mechanism in such models to make multimodal fusion more scalable with an increasing number of modalities.
+In summary, across diverse clinical datasets and modalities, OvO attention consistently outperforms traditional fusion techniques in both predictive performance and computational efficiency, underlining its robustness in handling complex multimodal healthcare data.
 6 Conclusion
@@
 The proposed formulation significantly reduces the computational complexity compared to the widely used early fusion through self-attention and cross-attention methods.
-Notably, OvO achieves, at minimum, a reduction of 91.98% in FLOPs when benchmarked against self and cross-attention methods across a range of clinical datasets encompassing four, five, and six modalities.
+Notably, OvO achieves, at minimum, a reduction of 91.98% in FLOPs when benchmarked against self and cross-attention methods across a range of clinical datasets containing up to six modalities.
 We provide both a detailed theoretical complexity analysis and empirical evidence from a simulated experiment, illustrating that OvO’s computational demand scales linearly with the number of modalities, in contrast to the quadratic scaling observed in other methods.
-Furthermore, the versatility of OvO is evidenced through its application to a non-clinical dataset, where it consistently surpassed all comparative baselines.
-Overall, the results unequivocally establish that OvO not only significantly reduces computational expenses but also meets or exceeds the performance of existing state-of-the-art fusion methodologies.
-8 Impact Statement
-Our proposed work, OvO, provides a method to overcome one of the major challenges associated with multimodal datasets - computational resource demand and cost.
-The scalability of the proposed formula allows diverse multimodal experiments to run in a wider range of research environments despite economic and computational constraints.
-While our method of modality fusion is meant to reduce the negative environmental impact of attention-learning models, we note that the individual modality encoders used (e.g., ClinicalBERT) still have a negative impact by creating emissions.
+Our proposed method provides a way to overcome one of the major challenges associated with multimodal datasets - computational resource demand and cost, thus enabling adoption in resource-constrained domains, such as clinical decision support.
+Overall, the results unequivocally establish that OvO not only significantly reduces computational expenses but also exceeds the performance of existing state-of-the-art fusion methodologies.
 Appendix A Computational Complexity Analysis for Multimodal Integration Schemes
@@
 Attention scores are computed by taking the dot product of queries and keys.
-The self-attention mechanism has quadratic complexity with respect to the sequence length and linear complexity with respect to the representation dimension [Vaswani et al., 2017].
+The self-attention mechanism has quadratic complexity with respect to the sequence length and linear complexity with respect to the representation dimension [32].
 Thus, given the concatenated sequence’s length of and the dimension of the keys and queries , the complexity of this step is:
@@
 Thus, we chose 0.15 as the threshold.
-Appendix C eICU Modality Descriptions and Detailed Pre-processing
-The patient table contains patient demographic, admission, and discharge details.
-We use this table to determine mortality status.
-The diagnosis table contains active diagnoses given to each patient.
-We extract diagnosis features by one-hot encoding conditions specified in the ‘diagnosisstring’ column.
-The treatment table contains active treatments prescribed for each patient.
-We extract treatment features by one-hot encoding treatment types specified in the ‘treatmentstring’ column.
-The medication contains active medication orders for patients and when they were ordered.
-We extract medication features by one-hot encoding drugs specified in the ‘drugname’ column.
-The lab table contains lab type and corresponding results for various lab measurements collected for each patient.
-We extract lab features by summing lab measurements of commonly recorded lab types for each patient stay.
-Appendix D TCGA Modality Descriptions and Detailed Pre-processing
-CNV defines the varying number of repeats of genetic fragments found in a human genome.
-The number of repeats of specific genetic fragments influences gene expression levels and has been associated with the progression of different cancers [Samulin Erdem et al., 2017].
-Any genomic regions missing CNV values or only having one unique value across all cases were removed.
-DNA methylation represents the amount of condensation of genetic regions due to the chemical alteration imposed by methyl groups.
-This condensation generally represses gene activity near the genetic region.
-Any genomic regions with missing values were removed.
-Clinical data includes information such as the patient’s diagnosis, demographics, laboratory tests, and family relationships.
-Categorical features were isolated and a coefficient of variation test was run to determine highly variable features.
-Features with a coefficient of variation higher than 70 were kept for analysis, along with the target variable.
-These features were converted into numerical format using one-hot-encoding.
-Gene expression data is collected through RNA-sequencing.
-Levels of gene expression are recorded by detecting the amounts of transcripts found for each gene.
-These levels can be used to determine the molecular mechanisms underlying cancer.
-Transcriptomic data was filtered to only include protein-coding genes and measured in fragments per kilobase of exon per million mapped fragments (FPKM).
-Imaging - TCGA collects pathology slide images of tissues sampled from the tumor.
-This modality provides visual information about the malignant region and can help with diagnosis and treatment planning.
-The image data was filtered only to include DX images, which result from a single X-Ray exposure, rotated to landscape view, then cropped to the median aspect ratio of 1.3565.
-We filtered for patients that had all five modalities, and we also only chose the patients that were still alive, to create a more balanced number of patients between cancer types (338 colon cancer patients, 329 kidney, 301 lung, 228 liver, and 226 stomach patients, after the filtering).
-The task we created is to classify each patient’s cancer type.
-For all modalities, features with missing values were dropped.
-For CNV, DNA Methylation, and gene expression data, feature reduction was performed using a random forest classifier, only on training data, ensuring the test was not seen by the random forest.
-Using the validation set, we determined the best number of estimators (out of 50, 100, 150).
-Appendix E Hyperparameter Tuning
-For each experiment, we used the validation accuracy to determine the best hyperparameters.
-We tuned the learning rate (0.01 - ), batch size (16, 32, 64, 128), epochs (200 epochs with early stopping if validation accuracy did not increase for 5 epochs), and number of attention heads for the OvO and pairwise cross-modal attention models (1, 2, 4, 8, 16).
-For the neural network encoders, we tuned the number of linear layers ranging from 1 to 4.
-Similarly, for the convolutional neural network, we tuned the number of convolution layers ranging from 1 to 4.
-Appendix F Amazon Reviews Dataset Description and Detailed Pre-processing
-The Amazon customer review dataset [Ni et al., 2019] aims to understand consumer opinions and experiences with products sold on Amazon.
-We sampled from the Electronics category, as it was one of the largest and most commonly used [Li et al., 2021, Haque et al., 2018, Pandey and Joshi, 2022].
-We used three modalities from the dataset – review images, text, and the metadata (price, brand name, product category, etc.) associated with each review.
-The task is to classify the binary sentiment of each review - positive (4-5 stars) or negative (1-2 stars), with a total of around 20,000 samples.
-While sentiment classification with the Amazon reviews data is well-studied unimodally through text [Haque et al., 2018, Li et al., 2021, 2023, Pandey and Joshi, 2022], we offer a unique multimodal approach to the task.
-Existing multimodal approaches on Amazon reviews focus on review helpfulness [Han et al., 2022, Liu et al., 2021], or product similarity [Das et al., 2022] rather than sentiment classification.
-Furthermore, other papers use only text and images or only images and metadata (as text); an image-text-tabular multimodal task has yet to be explored.
-Intuitively, images and metadata can both be useful for review sentiment prediction - if the image highlights a broken item with the metadata showing a high price or trusted brand, consumer dissatisfaction may be implied.
-Appendix G Compute Resources
+Appendix C Compute Resources
 For each experiment, we use one NVIDIA GeForce RTX 3090 GPU.
-For the Hateful Memes task, single-modality models ran for roughly 40 minutes, and multi-modal models ran for roughly 55 minutes on average.
-For the Amazon reviews task, the single modality pre-trained models ran for roughly 50 minutes, the single modality neural network ran for a minute, and the multi-modal models ran for approximately an hour on average.
-For the TCGA task, single-modality models ran for 5 minutes, while multi-modal models ran for roughly 15 minutes on average.
+For the MIMIC task, single-modality models ran for roughly 40 minutes, and multi-modal models ran for roughly 55 minutes on average.
+For the eICU, the single modality pre-trained models ran for roughly 50 minutes, the single modality neural network ran for a minute, and the multi-modal models ran for approximately an hour on average.
+For the TADPOLE task, single-modality models ran for 5 minutes, while multi-modal models ran for roughly 15 minutes on average.
 In the simulation dataset, the maximum modalities was 20 which took our model, OvO, roughly 2 minutes to run, while the cross-modal attention baseline took about 20 minutes to run on average.
-Appendix H Significance Testing
-We use a t-test to determine if there is a significant difference the performance metrics (F1-score, accuracy, AUROC, AUPRC, MAUC, BCA) means between OvO attention and the next best-performing multimodal model.
+[…]
+Appendix D Significance Testing
+We use a t-test to determine if there is a significant difference the performance metrics (AUROC, AUPRC, MAUC, BCA) means between OvO attention and the next best-performing multimodal model.
 Our sample size is 10 from each group, as we initialized the models with 10 random seeds.
@@
 The p-value for the AUROC scores is 0.00363 and the p-value for AUPRC is 0.000948.
-For the TCGA dataset, we do not have evidence to reject the null hypothesis and cannot say that the accuracy and F1-score means were different between OvO and cross-attention since the p-values were greater than = 0.01 (p-value of 0.04 for accuracy means, and p-value of 0.02 for F1-score means).
-This demonstrates that although cross-attention performed slightly better than OvO, it was not statistically significant.
 For the TADPOLE challenge, we compare against cross-attention as it performed the second best after OvO.
 We get a p-value for MAUC scores of and a p-value of for BCA.
-Thus, we demonstrate a statistically significant difference in accuracy and F1-score means between self-attention and OvO attention.
+Thus, we demonstrate a statistically significant difference in MAUC and BCA means between self-attention and OvO attention.
 Lastly, for the eICU dataset, we compare against cross-attention as it performed the second best after OvO.
 We get a p-value for AUROC scores of and a p-value of for AUPRC.
-Thus, we demonstrate a statistically significant difference in accuracy and F1-score means between self-attention and OvO attention.
+Thus, we demonstrate a statistically significant difference in AUROC and AUPRC means between self-attention and OvO attention.
+[…]
+Appendix E Limitations
+This paper’s primary goal is to address one of the major challenges associated with multimodal datasets - computational resource demand and cost.
+While we demonstrate the scalability of the OvO attention mechanism and its efficiency in handling multiple modalities, we did not conduct experiments focused on interpretability, which is also a crucial aspect of multimodal learning in the biomedical space.
+The potential for OvO to reveal modality importance through the learninable parameter is promising, but further experiments are required to explore this interpretability in a meaningful way.
+Additionally, while we compared OvO to popular attention mechanisms like self-attention and cross-attention, there are other attention variants that could serve as future baselines, though they are not widely used in multimodal models.
+Lastly, the publicly available multimodal datasets often limit the diversity of modalities we can test, with many datasets primarily consisting of tabular data.